# Probing Proton versus Electron Heating and Energization during Magnetic Reconnection

### v1

@@ 1 Introduction
Magnetic reconnection is responsible for the rapid conversion of magnetic energy in various environments, including solar flares (R. P. Lin & H. S. Hudson, 1971; S. Masuda et al., 1994; R. C. Lin et al., 2003), Earth’s magnetosphere (J. W. Dungey, 1961; B. U. Ö. Sonnerup et al., 1981), and the solar wind (J. T. Gosling et al., 2005; T. D. Phan et al., 2006; T. Phan et al., 2021). During reconnection, magnetic energy is efficiently transferred to particles, producing both bulk heating that scales with $m_{i}C_{A}^{2}$ (the available magnetic energy per particle) (T. D. Phan et al., 2013, 2014; M. Øieroset et al., 2023, 2024; M. Oka et al., 2025) and a significant number of non-thermal particles characterized by power-law tails in their distribution functions (R. C. Lin et al., 2003; M. Øieroset et al., 2002; S. Krucker et al., 2010; D. E. Gary et al., 2018; R. Ergun et al., 2020a; M. I. Desai et al., 2025). In situ measurements of bulk heating have revealed that protons gain significantly more energy than electrons and there is some evidence that protons may also dominate energy gain in the non-thermal component during magnetotail reconnection (R. Ergun et al., 2020b; A. Rajhans et al., 2025). In the case of solar flares, measurements of energetic protons do not extend below around an MeV so the total energy content of protons is uncertain (R. C. Lin et al., 2003; A. Emslie et al., 2012).
@@
Magnetic reconnection produces bent magnetic field lines that expand outward due to their tension force to form an Alfvénic exhaust (E. N. Parker, 1957; H. E. Petschek, 1964; T. Sato & T. Hayashi, 1979). The exhaust carries energy away from the x-line and transfers energy from the magnetic field to the surrounding plasma (Y. Lin & L. C. Lee, 1993). Reconnecting current sheets, however, tend to breakup and form multiple x-lines and associated flux ropes (D. Biskamp, 1986; J. F. Drake et al., 2006b; N. Loureiro et al., 2007; A. Bhattacharjee et al., 2009; W. Daughton et al., 2011) and observations support the multi-x-line picture of reconnection (L.-J. Chen et al., 2008; T. Phan et al., 2024) The dominant mechanism driving particle energy gain during magnetic reconnection is Fermi reflection in growing and merging magnetic flux ropes (J. F. Drake et al., 2006b; M. Oka et al., 2010; J. Dahlin et al., 2014; F. Guo et al., 2014; X. Li et al., 2019; Q. Zhang et al., 2021). The mechanism produces power-law tails in both electron (H. Arnold et al., 2021) and ion (Q. Zhang et al., 2021; Z. Yin et al., 2024a) distributions. Particles gain energy through repeated reflections in contracting magnetic field lines, thereby leading to the observed power-law tails in particle energy distributions (J. F. Drake et al., 2006b; J. Drake et al., 2013; X. Li et al., 2019; Q. Zhang et al., 2021). The rate of energy gain from Fermi reflection is proportional to a particle’s energy (J. F. Drake et al., 2006b; J. Drake et al., 2013) and is therefore greatest for the most energetic particles. For this reason, the most energetic particles gain the most energy, which facilitates the formation of extended powerlaw tails. The kglobal model (H. Arnold et al., 2019; J. F. Drake et al., 2019; Z. Yin et al., 2024b), which is designed to describe reconnection in macroscale systems by ordering out all kinetic scales, was the first fully self-consistent model to produce the extended powerlaw tails documented in observations (H. Arnold et al., 2019; J. F. Drake et al., 2019).
@@
The kglobal model was recently upgraded to include particle protons (Z. Yin et al., 2024b) and simulations using this model revealed that the energy content of energetic protons exceeded that of electrons (Z. Yin et al., 2024a) even when both species start with equal initial temperatures. Earlier particle-in-cell (PIC) simulations also explored the relative heating of electrons and protons during reconnection (C. C. Haggerty et al., 2015). These simulations established that the proton temperature gain exceeds that of the electrons during magnetic reconnection. However, the underlying mechanisms that facilitate proton versus electron heating and energization have not been identified and the PIC simulations were unable to match the observed non-thermal spectra of the two species.
@@ 2 Simulation Model Setup
Grid Points 2048×10242048$\times$ 10242048 × 1024
@@ 3 Simulation Results
Figure 2: Time evolution of the energy spectra of electrons (dashed lines) and protons (solid lines) during the simulation with equal initial temperatures. The initial spectrum (proton/electron)is shown as a black solid line. Subsequent spectra are plotted at $t/\tau_{A}=2$, 3, 6, and 18.
@@
To further confirm that the initial energy gain of electrons is controlled by Fermi reflection on entry into reconnection exhausts, we explore the mass dependence of early time electron heating with simulations employing varying electron-to-proton masses but with fixed upstream electron temperatures. In this situation, Eq. (2) predicts that the initial electron energy gain should scale as $(m_{e}/m_{i})^{1/2}$. Shown in Figure 5 are cuts along the center of the current sheet of the electron parallel temperature at $t=2.2\,\tau_{A}$ from three simulations, differing only by their mass ratio: $m_{i}/m_{e}=25$, 100, and 400. The ion mass $m_{i}$ is fixed in all cases, so a higher mass ratio corresponds to a lower electron mass. At $t=2.2\,\tau_{A}$ the electrons have interacted with a single reconnection outflow (see discussion related to Figure 4). The cuts reveal that the simulation with $m_{i}/m_{e}=25$ (heavier electrons) produces the strongest electron energy gain, while the case with $m_{i}/m_{e}=400$ (i.e., lighter electrons) produces the least. To quantify this result, we have calculated the average parallel electron temperature increment within the outmost magnetic separatrix at $t=2.2\tau_{A}$ for each of the three simulations. The average temperature increments are 0.065, 0.045, and 0.038 for $m_{i}/m_{e}=25$, 100100100 and 400400400, respectively. Thus, this early time temperature increment decreases with electron mass but not as much as predicted by the $(m_{e}/m_{i})^{1/2}$ scaling of Eq. (2). This is likely because, even at this early time, the high velocity electrons undergo many bounces in small magnetic flux ropes. This is especially true for the lower-electron-mass cases where the initial thermal speed greatly exceeds the Alfvén speed. In any case, the increased early heating of protons compared with electrons is a consequence of greater energy gain during Fermi reflection in developing reconnection exhausts.
@@
Thus, the simulations suggest that protons gain substantially more energy than electrons during the early phase of reconnection and that their energy gain never overcomes that of the protons. That is, since the dominant mechanism for energy gain is Fermi reflection during flux rope merger, in which the rate of energy gain is proportional to energy (J. F. Drake et al., 2006a; J. Drake et al., 2013), once the proton energy exceeds that of electrons at early time the protons continue to gain energy faster and thus, reach higher energies over the full range of energies probed in the simulations.
@@
Figure 6: Energy spectra of electrons and protons from four simulations with different initial electron-to-proton temperature ratios. The initial proton temperature is held constant, while the initial electron temperature is varied as $T_{e(\mathrm{initial})}=0.5T_{i(\mathrm{initial})}$, $T_{e(\mathrm{initial})}=T_{i(\mathrm{initial})}$, $T_{e(\mathrm{initial})}=2T_{i(\mathrm{initial})}$, and $T_{e(\mathrm{initial})}=5T_{i(\mathrm{initial})}$, as indicated in each panel. The energy spectra are computed at late time, after the formation of a dominant magnetic island in the current sheet, by summing particles across the entire simulation domain. Dashed lines represent initial spectra, while solid lines indicate late-time spectra for protons (blue) and electrons (orange).
@@ 4 Conclusion
In situ observations of electron and proton heating during reconnection in near-Earth space have revealed that heating of both species scales with the available magnetic energy per particle $m_{i}C_{A}^{2}$, with protons gaining significantly more energy than electrons ($0.17\,m_{i}C_{A}^{2}$ for protons and $0.02\,m_{i}C_{A}^{2}$ for electrons (T. D. Phan et al., 2013, 2014; M. Øieroset et al., 2023, 2024), although in the magnetotail observations there was some reduction of proton heating below this scaling for large values of $C_{A}$ (M. Øieroset et al., 2024)). This result was interpreted as arising from the electric potential that develops between the center of the reconnecting current sheet and the upstream, which can be substantial once the low density lobe plasma participates in the reconnection. PIC simulations of electron and proton heating during reconnection reproduce this basic trend (M. A. Shay et al., 2014; C. C. Haggerty et al., 2015). However, an understanding of why proton heating is much stronger than electron heating remains elusive. More recently, measurements have been extended to much higher energies, which are characterized by power-law distributions. There is some evidence that the power-law tails of protons contain more energy than those of electrons (R. Ergun et al., 2020a) and the first comprehensive study establishing this has recently been completed (A. Rajhans et al., 2025). Further, in flare observations, where in situ measurements are not possible, the measurement of protons below an MeV are not available, so estimates of the relative energy content of the non-thermal components of the two species are uncertain (A. Emslie et al., 2012). Recent simulations of electron and proton energy gain during reconnection, however, have revealed that non-thermal protons gain more energy than non-thermal electrons over several decades of energy (Z. Yin et al., 2024a). In the absence of observational support for this result during flare energy release, it is essential to establish the underlying physical mechanism responsible.
In this paper, we have investigated proton and electron heating and energization during magnetic reconnection using the kglobal simulation model. This is presently the only model that is able to produce the extended powerlaw energy spectra seen in observations. The simulations reveal that the protons experience a strong jump in energy early in time as they first enter into reconnection exhausts (Figures 1 and 7). This energy jump greatly exceeds that of electrons. Since energy gain as time progresses is dominated by Fermi reflection in merging flux ropes, where energy gain is proportional to particle energy (J. F. Drake et al., 2006b; M. Oka et al., 2010; J. Drake et al., 2013), the initial strong heating of protons facilitates their increased energy gain compared with electrons for the duration of the reconnection dynamics.
@@
Beyond the overall energy spectrum of electrons and protons, we have focused on obtaining an explicit value for the electron and proton increment during reconnection. We emphasize that since the energy normalization of the kglobal model is $m_{i}C_{A}^{2}$ the scaling of heating of both species scales with this parameter, as documented in observations. Since the value of the electron mass in the simulations is artificial, the sensitivity of the result to $m_{i}/m_{e}$ was checked with simulations with mass ratios $m_{i}/m_{e}=25$, 100100100 and 400400400. The electron and proton temperature increments for simulations with these three mass-ratios are, respectively: $\Delta T_{e}=0.070$, 0.0790.0790.079 and 0.0710.0710.071; and $\Delta T_{i}=0.313$, 0.3060.3060.306 and 0.3140.3140.314. Thus, the late time relative temperature increments of the two species are insensitive to the mass-ratio, a result that is consistent with previous PIC simulations (M. A. Shay et al., 2014). This result would seem to be at odds with the conclusions of M. Oka et al. (2025) who concluded that the ratio of proton to electron heating scales as $(m_{p}/m_{e})^{1/4}$ based on the analysis of in situ observations. However, the mass ratio in the observations was clearly 1836 and was not varied. Thus, their scaling yields a temperature ratio of around 6.5 and no conclusion about the scaling with $m_{p}/m_{e}$ was possible from this data. Finally, we emphasize that the increased energy increment of protons compared to electrons persists at high energy in the non-thermal spectra of the two species.
@@
A recent survey of energetic reconnection events in the Earth’s magnetotail provides important observational constraints on eletron and proton heating and acceleration (A. Rajhans et al., 2025). Consistent with the present simulations, these observations revealed that protons consistently gain more energy than electrons across the full range of particle energies (in the thermal and non-thermal components of both species). However, there are differences between the observations and modeling results. The observations revealed that electrons typically have harder spectra even though the non-thermal electrons carried less energy than the non-thermal protons. In the present simulations the spectral indices of the non-thermal particles are comparable (see Fig. 6) even when the initial ratio of electron and proton temperatures varies. The electron spectra at high energy are slightly harder than those of the protons. A clear difference between the simulations and the observations is the measured (in observations) inverse correlation between the fraction of non-thermal particles and the measured temperature of those particles. In the present model, the non-thermal fraction of both species depends strongly on the strength of the guide field and not the value of $m_{i}C_{A}^{2}$ (which controls the temperature) (H. Arnold et al., 2021; Z. Yin et al., 2024a). On the other hand, the present model does not include some potentially important physics. In the simulations there is no mechanism for energetic particles to escape from the energy release region, a process that could limit the energy content of the non-thermal particles. Finally, the demagnetization of protons at high energy, which is also not included in the kglobal model, might limit the energy gain of the most energetic non-thermal protons. Further exploration of these topics in the models is warranted.
@@
We acknowledge extensive discussions with Drs. Tai Phan, Marit Oieroset, Mitsuo Oka and the MMS Science Team. Support was provided from NASA Grant Nos. 80NSSC20K1277, 80NSSC20K1813 and 80NSSC22K0352, and NSF Grant No. PHY2109083. The simulations were carried out at the National Energy Research Scientific Computing Center (NERSC). The data used to perform the analysis and construct the figures for this paper are preserved at the NERSC High Performance Storage System and are available upon request.

### v2

@@ 1 Introduction
Magnetic reconnection is responsible for the rapid conversion of magnetic energy in various environments, including solar flares (lin71; Masuda94; Lin03; Benz17), Earth’s magnetosphere (Dungey61; Sonnerup81), and the solar wind (Gosling05; Phan06; phan21). During reconnection, magnetic energy is efficiently transferred to particles, producing both bulk heating that scales with $m_{i}C_{A}^{2}$ (the available magnetic energy per particle) (Phan13a; Phan14; Oieroset23; Oieroset24; Oka25) and a significant number of non-thermal particles characterized by power-law tails in their distribution functions (Lin03; Oieroset02; Krucker10; Gary18; Ergun20b; Desai25). In situ measurements of bulk heating have revealed that protons gain significantly more energy than electrons and there is some evidence that protons may also dominate energy gain in the non-thermal component during magnetotail reconnection (Ergun20; Rajhans25). In the case of solar flares, measurements of energetic protons do not extend below around an MeV so the total energy content of protons is uncertain (Lin03; Emslie12). However, recent observations of line broadening in flares suggests ion heating in flares might also exceed that of electrons (Russell25).
@@
Magnetic reconnection produces bent magnetic field lines that expand outward due to their tension force to form an Alfvénic exhaust (Parker57; Petschek64; Sato79). The exhaust carries energy away from the x-line and transfers energy from the magnetic field to the surrounding plasma (Lin93). Reconnecting current sheets, however, tend to breakup and form multiple x-lines and associated flux ropes (Biskamp86; Drake06; Loureiro07; Bhattacharjee09; Daughton11) and observations support the multi-x-line picture of reconnection (Chen08; Phan24) The dominant mechanism driving particle energy gain during magnetic reconnection is Fermi reflection in growing and merging magnetic flux ropes (Drake06; Oka10; Dahlin14; Guo14; Li19; Zhang21). The mechanism produces power-law tails in both electron (Arnold21) and ion (Zhang21; Yin24b) distributions. Particles gain energy through repeated reflections in contracting magnetic field lines, thereby leading to the observed power-law tails in particle energy distributions (Drake06; Drake13; Li19; Zhang21). The rate of energy gain from Fermi reflection is proportional to a particle’s energy (Drake06; Drake13) and is therefore greatest for the most energetic particles. For this reason, the most energetic particles gain the most energy, which facilitates the formation of extended powerlaw tails. The kglobal model (Arnold19; Drake19; Yin24), which is designed to describe reconnection in macroscale systems by ordering out all kinetic scales, was the first fully self-consistent model to produce the extended powerlaw tails documented in observations (Arnold19; Drake19).
@@
The kglobal model was recently upgraded to include particle protons (Yin24) and simulations using this model revealed that the energy content of energetic protons exceeded that of electrons (Yin24b) even when both species start with equal initial temperatures. Earlier particle-in-cell (PIC) simulations also explored the relative heating of electrons and protons during reconnection (Haggerty15). These simulations established that the proton temperature gain exceeds that of the electrons during magnetic reconnection. However, the underlying mechanisms that facilitate proton versus electron heating and energization have not been identified and the PIC simulations were unable to match the observed non-thermal spectra of the two species.
@@ 2 Simulation Model Setup
Grid Points $2048\times 1024$
@@ 3 Simulation Results
Figure 2: Time evolution of the energy spectra of electrons (dashed lines) and protons (solid lines) during the simulation with equal initial temperatures. The initial spectrum (proton/electron)is shown as a black solid line. Subsequent spectra are plotted at $t/\tau_{A}=2$, 3, 6, and 18. A black dashed line line with a slope of -3.6 matches the slope of the late-time proton spectrum. The late time powerlaw index of the electrons is around -3.2.
@@
To further confirm that the initial energy gain of electrons is controlled by Fermi reflection on entry into reconnection exhausts, we explore the mass dependence of early time electron heating with simulations employing varying electron-to-proton masses but with fixed upstream electron temperatures. In this situation, Eq. (2) predicts that the initial electron energy gain should scale as $(m_{e}/m_{i})^{1/2}$. Shown in Figure 5 are cuts along the center of the current sheet of the electron parallel temperature at $t=2.2\,\tau_{A}$ from three simulations, differing only by their mass ratio: $m_{i}/m_{e}=25$, 100, and 400. The ion mass $m_{i}$ is fixed in all cases, so a higher mass ratio corresponds to a lower electron mass. At $t=2.2\,\tau_{A}$ the electrons have interacted with a single reconnection outflow (see discussion related to Figure 4). The cuts reveal that the simulation with $m_{i}/m_{e}=25$ (heavier electrons) produces the strongest electron energy gain, while the case with $m_{i}/m_{e}=400$ (i.e., lighter electrons) produces the least. To quantify this result, we have calculated the average parallel electron temperature increment within the outmost magnetic separatrix at $t=2.2\tau_{A}$ for each of the three simulations. The average temperature increments are 0.065, 0.045, and 0.038 for $m_{i}/m_{e}=25$, $100$ and $400$, respectively. Thus, this early time temperature increment decreases with electron mass but not as much as predicted by the $(m_{e}/m_{i})^{1/2}$ scaling of Eq. (2). This is likely because, even at this early time, the high velocity electrons undergo many bounces in small magnetic flux ropes. This is especially true for the lower-electron-mass cases where the initial thermal speed greatly exceeds the Alfvén speed. In any case, the increased early heating of protons compared with electrons is a consequence of greater energy gain during Fermi reflection in developing reconnection exhausts.
@@
Thus, the simulations suggest that protons gain substantially more energy than electrons during the early phase of reconnection and that their energy gain never overcomes that of the protons. That is, since the dominant mechanism for energy gain is Fermi reflection during flux rope merger, in which the rate of energy gain is proportional to energy (Drake06a; Drake13), once the proton energy exceeds that of electrons at early time the protons continue to gain energy faster and thus, reach higher energies over the full range of energies probed in the simulations.
@@
Figure 6: Energy spectra of electrons and protons from four simulations with different initial electron-to-proton temperature ratios. The initial proton temperature is held constant, while the initial electron temperature is varied as $T_{e(\mathrm{initial})}=0.5T_{i(\mathrm{initial})}$, $T_{e(\mathrm{initial})}=T_{i(\mathrm{initial})}$, $T_{e(\mathrm{initial})}=2T_{i(\mathrm{initial})}$, and $T_{e(\mathrm{initial})}=5T_{i(\mathrm{initial})}$, as indicated in each panel. The energy spectra are computed at late time, after the formation of a dominant magnetic island in the current sheet, by summing particles across the entire simulation domain. Dashed lines represent initial spectra, while solid lines indicate late-time spectra for protons (blue) and electrons (orange). At late time the spectral indices with the various initial electron temperatures differ only slightly from the equal temperature simulation. The powerlaw slopes of the proton spectra in these settings range from −4.0 to −3.6, whereas those of the electron spectra range from −3.2 to −3.0
@@ 4 Conclusion
In situ observations of electron and proton heating during reconnection in near-Earth space have revealed that heating of both species scales with the available magnetic energy per particle $m_{i}C_{A}^{2}$, with protons gaining significantly more energy than electrons ($0.17\,m_{i}C_{A}^{2}$ for protons and $0.02\,m_{i}C_{A}^{2}$ for electrons (Phan13a; Phan14; Oieroset23; Oieroset24), although in the magnetotail observations there was some reduction of proton heating below this scaling for large values of $C_{A}$ (Oieroset24)). This result was interpreted as arising from the electric potential that develops between the center of the reconnecting current sheet and the upstream, which can be substantial once the low density lobe plasma participates in the reconnection. PIC simulations of electron and proton heating during reconnection reproduce this basic trend (Shay14; Haggerty15). However, an understanding of why proton heating is much stronger than electron heating remains elusive. More recently, measurements have been extended to much higher energies, which are characterized by power-law distributions. There is some evidence that the power-law tails of protons contain more energy than those of electrons (Ergun20b) and the first comprehensive study establishing this has recently been completed (Rajhans25). Further, in flare observations, where in situ measurements are not possible, the measurement of protons below an MeV are not available, so estimates of the relative energy content of the non-thermal components of the two species are uncertain (Emslie12). Recent simulations of electron and proton energy gain during reconnection, however, have revealed that non-thermal protons gain more energy than non-thermal electrons over several decades of energy (Yin24b). In the absence of observational support for this result during flare energy release, it is essential to establish the underlying physical mechanism responsible.
In this paper, we have investigated proton and electron heating and energization during magnetic reconnection using the kglobal simulation model. This is presently the only model that is able to produce the extended powerlaw energy spectra seen in observations. The simulations reveal that the protons experience a strong jump in energy early in time as they first enter into reconnection exhausts (Figures 1 and 7). This energy jump greatly exceeds that of electrons. Since energy gain as time progresses is dominated by Fermi reflection in merging flux ropes, where energy gain is proportional to particle energy (Drake06; Oka10; Drake13), the initial strong heating of protons facilitates their increased energy gain compared with electrons for the duration of the reconnection dynamics.
@@
Beyond the overall energy spectrum of electrons and protons, we have focused on obtaining an explicit value for the electron and proton increment during reconnection. We emphasize that since the energy normalization of the kglobal model is $m_{i}C_{A}^{2}$ the scaling of heating of both species scales with this parameter, as documented in observations. Since the value of the electron mass in the simulations is artificial, the sensitivity of the result to $m_{i}/m_{e}$ was checked with simulations with mass ratios $m_{i}/m_{e}=25$, $100$ and $400$. The electron and proton temperature increments for simulations with these three mass-ratios are, respectively: $\Delta T_{e}=0.070$, $0.079$ and $0.071$; and $\Delta T_{i}=0.313$, $0.306$ and $0.314$. Thus, the late time relative temperature increments of the two species are insensitive to the mass-ratio, a result that is consistent with previous PIC simulations (Shay14). This result would seem to be at odds with the conclusions of Oka25 who concluded that the ratio of proton to electron heating scales as $(m_{p}/m_{e})^{1/4}$ based on the analysis of in situ observations. However, the mass ratio in the observations was clearly 1836 and was not varied. Thus, the observation of a temperature ratio of around 6.5 is an empirical result, rather than the result of a scaling with $m_{p}/m_{e}$. Finally, we emphasize that the increased energy increment of protons compared to electrons persists at high energy in the non-thermal spectra of the two species.
@@
A recent survey of energetic reconnection events in the Earth’s magnetotail provides important observational constraints on eletron and proton heating and acceleration (Rajhans25). Consistent with the present simulations, these observations revealed that protons consistently gain more energy than electrons across the full range of particle energies (in the thermal and non-thermal components of both species). However, there are differences between the observations and modeling results. The observations revealed that electrons typically have harder spectra even though the non-thermal electrons carried less energy than the non-thermal protons. In the present simulations the spectral indices of the non-thermal particles are comparable (see Fig. 6) even when the initial ratio of electron and proton temperatures varies. The electron spectra at high energy are slightly harder than those of the protons. A clear difference between the simulations and the observations is the measured (in observations) inverse correlation between the fraction of non-thermal particles and the measured temperature of those particles. In the present model, the non-thermal fraction of both species depends strongly on the strength of the guide field and not the value of $m_{i}C_{A}^{2}$ (which controls the temperature) (Arnold21; Yin24b). On the other hand, the present model does not include some potentially important physics. In the simulations there is no mechanism for energetic particles to escape from the energy release region, a process that could limit the energy content of the non-thermal particles. Finally, the demagnetization of protons at high energy, which is also not included in the kglobal model, might limit the energy gain of the most energetic non-thermal protons. Further exploration of these topics in the models is warranted.
The simulations presented in this manuscript are based on the assumption that classical collisions can be neglected. In the case of the solar wind and the Earth space environment such an assumption is reasonable because of the very low plasma densities and associated weak classical collisions. In the case of solar flares, however, the neglect of classical collisions is not as obvious. A key parameter is the strength of the reconnection electric field compared with the Dreicer runaway electric field $E_{D}$, which is the electric field needed to accelerate electrons to their thermal speed during a classical collision time (Dreicer60),
$eE_{D}=m_{e}\nu_{e}v_{the},$ (3)
with $v_{the}$ the electron thermal speed and $\nu_{e}$ the electron collision rate,
$\nu_{e}=2.9\times 10^{-6}n_{e}\ln\Lambda/T_{e}^{3/2},$ (4)
where $\ln\Lambda$ is the Coulomb logarithm, the density $n_{e}$ is in units of $cm^{-3}$ and the electron temperature is in $eV$. For typical coronal ambient parameters ($n_{e}\sim 10^{10}/cm^{3}$ and $T_{e}\sim 100eV$), $\nu_{e}\sim 440/s$ and $E_{D}\sim 1.05\times 10^{-2}V/m$. In the collisionless limit the reconnection electric field $E_{r}$ is around $0.1C_{A}B$, where $C_{A}$ is the upstream Alfvén speed (Shay99; Shay07; Torbert18; Burch20). For a magnetic field of 50GG with the same density, $C_{A}\sim 1.1\times 10^{6}m/s$ and $E_{r}\sim 550V/m$, which is far above the Dreicer value. Thus, reconnection in solar flares is deeply in the collisionless regime.
On the other hand, the dominant driver of the energization of the highest energy electrons and ions in flares is the merger of flux ropes. The merger of two equal-sized flux ropes causes the energy of all particles within the flux ropes to approximately double (the field line shortens by a factor of $\sqrt{2}$ and, as a result of the second adiabatic invariant, the particle parallel velocity increases by the same factor) (Drake13). Classical collisional energy losses can be neglected if the merging time $\tau_{r}$ is less than $\nu_{e}^{-1}$. The merging time of two flux ropes of radii $R$ is therefore given by $10R/C_{A}$ and the neglect of classical energy loss requires
$\nu_{e}<0.1C_{A}/R.$ (5)
Reconnection should onset with the smallest flux ropes that can magnetize protons or $R\sim 5d_{i}$ with $d_{i}=c/\omega_{pi}=C_{A}/\Omega_{ci}$ the ion inertial length scale (Mandt94). Thus, the neglect of collisional losses requires
$\nu_{e}<0.02\Omega_{ci}$ (6)
with $\Omega_{ci}$ the proton cyclotron frequency. This inequality is easily satisfied for the typical solar parameters previously discussed. Further, the inequality in Equation (5) will continue to be satisfied during subsequent island mergers. The collision rate scales as $v^{-3}$. Since $v$ increases by the factor $\sqrt{2}$ during each merger, after $N$ mergers $\nu_{e}$ is reduced by $2^{-3N/2}$. The merging time $\tau_{r}$, however, increases with each merger: $\tau_{r}=10R/C_{A}\propto R/B\sim R^{2}\sim 2^{N}$ since the magnetic flux $BR$ is invariant during island merger (Fermo10). Thus, $\nu_{e}\tau_{r}\propto 2^{-N/2}$. If Equation (5) is satisfied, $\nu_{e}\tau_{r}<1$ for all subsequent mergers as particles gain energy.
Also of interest in solar flares is the rate at which particles gain energy and whether it is fast enough to be consistent with the rise time of hard x-rays. The shortest time scales are likely to be associated with island merger. The merger time depends, of course, on its radius as well as the local upstream Alfvén speed. Since a range of island sizes are expected in a reconnecting current layer (Fermo10), there is no universal time associated with island merger. For an island with scale size around 100kmkm with $C_{A}\sim 1000km/s$ as discussed above, the merger time and therefore the energy doubling time is around one second. Smaller islands can merge more quickly and are also able to drive strong particle energy gain since even the Larmor radii of protons in the MeV range are only hundreds of meters and are therefore well magnetized.
@@
We acknowledge extensive discussions with Drs. Tai Phan, Marit Øieroset, Mitsuo Oka and the MMS Science Team. Support was provided from NASA Grant Nos. 80NSSC20K1277, 80NSSC20K1813 and 80NSSC22K0352, and NSF Grant No. PHY2109083. The simulations were carried out at the National Energy Research Scientific Computing Center (NERSC). The data used to perform the analysis and construct the figures for this paper are preserved at the NERSC High Performance Storage System and are available upon request.